The Red Wings hotel has two revenue generating departments, which are rooms and F&B outlets. In 2010, rooms department had $120,000 in revenues and F&B department had $80,000 in revenues. In 2011, forecasts demonstrate that total revenue generated from the rooms department is 12.00% higher than the one in 2010. Also, there is a 4.00% increase in F&B revenues over the one in 2010. Based on the information given, what is the projected percentage change in total revenues in 2011 over 2010?
Answer #1
2010 Revenues 200000 =120000+80000
2011 Projected Revenues:
Rooms department 134400 =120000*1.12
F&B department 83200 =80000*1.04
2011 Projected Revenues 217600
Change in Total revenues 17600 =217600-200000
Divide by 2010 Revenues 200000
Projected percentage change in total revenues 8.80% increase